OpenStudy (anonymous):

What is the sum of the arithmetic sequence 6, 14, 22 …, if there are 26 terms?

OpenStudy (anonymous):

a=6 ,d=8 ,n=26 apply the rule Sn=n/2(2a+(n-1)d) _____ Where a= first term ,d=difference between eveery two terms ,n = number of all terms

OpenStudy (anonymous):

s26= 26/2 (2(6)+(26-1)8

OpenStudy (anonymous):

so the answer is 2756?

OpenStudy (anonymous):

right ":D

OpenStudy (anonymous):

What is the 25th term of the arithmetic sequence 9, 15, 21, … ? so with this one it would be s25=25/2 (2(9) + (25-1)6

OpenStudy (anonymous):

so the 25th term would be 153?

OpenStudy (anonymous):

yea,u calculate it XD